# Next Generation Equation-Free Multiscale Modelling of Crowd Dynamics via Machine Learning

### v2

@@ 4 Numerical results
Here, we present the results of the proposed framework on predicting pedestrian flow following the configuration of a rectangular corridor with an obstacle described in Section 3.1. As already discussed in Section 3.2, we first collected pedestrian trajectory data using the SFM microscopic agent-based simulator under a wide range of initial conditions. The resulting training dataset $\mathcal{X}_{tr}$ is used to train our framework, while the testing dataset $\mathcal{X}_{ts}$ is used to assess its efficiency and prediction accuracy.
In the first step, we extracted the normalized density fields data $X_{tr},X_{ts}$ from the discrete pedestrian positions in $\mathcal{X}_{tr},\mathcal{X}_{ts}$ using KDE, as described in Section 3.3. Next, we employed POD on $\mathcal{X}_{tr}$ to discover the latent space dimension; see Section 3.4. Following the criterion in Eq. (32), the minimum latent dimension required for retaining more than 99% of the energy in the training dataset is $d=13$ modes, as shown in Fig. 3(a).
@@
Figure 3: Accuracy of the restriction and lifting operators in Eqs. (20) and (21). Panel (a) shows the minimum number of POD modes (colored diamonds) required for retaining the desired percentage of energy in the training data set. Panel (b) depicts the average (over the $C=20$ cases with different initial conditions) reconstruction error $e^{2,rec}_{k}$ in Eq. (33) with $d=13$ POD modes, in time for the training and testing datasets.
Using the discovered POD basis $U_{d}$, we next constructed the restriction and lifting operators in Eqs. (20) and (21). To assess their accuracy, we computed the relative $L_{2}$ reconstruction error $e^{2,rec}_{k}$ in Eq. (33) across the $k=1,\ldots,K$ time steps of each trajectory. Fig. 3(b) displays the average relative error over all the $C=20$ cases with different initial conditions included in either the training or testing datasets. While initially the reconstruction error is $\sim$15%, it decreases to 10% at the first 30s30\ $\text${s} and further reduces to $\sim$7% after 50s50\ $\text${s} for both seen (training set) and unseen (testing set) data, indicating high accuracy of the restriction and lifting operators with a significantly reduced latent space dimension $d=13$. We note that this level of reconstruction error serves as an infimum approximation error baseline, i.e., the best possible error for the proposed framework at the ambient–density profiles–space. ROM forecasts in the long term are therefore expected to surpass this baseline error due to the additional error contributions from model-form error and training/generalization errors.
We next train MVAR and LSTM models to learn ROMs for the latent dynamics (see Section 2.2.3), where the latent variables $\boldsymbol{y}(t)\in\mathbb{R}^{13}$ are obtained from the density profiles of the training set via the restriction operator. To determine the lag window, we employed the BIC and AIC criteria described in Appendix D; BIC indicated a lag window of $w=4$, whereas AIC favored a wider (as expected) $w=9$. Both windows were used for training both MVAR and LSTM models, denoted from now on as MVAR(4), MVAR(9), LSTM(4), and LSTM(9) models. The training data sets were divided into features $Y_{f}=\{\boldsymbol{y}_{c}(t_{k-w}),\ldots,\boldsymbol{y}_{c}(t_{k-1})\}$ and targets $Y_{t}=\{\boldsymbol{y}_{c}(t_{k})\}$.
@@
Table 1: Training results of MVAR(4), MVAR(9), LSTM(4) and LSTM(9) models at the latent space. Loss functions $\mathcal{L}_{\text{MVAR}}\left(\cdot\right)$ in Eq. (25) for MVARs and $\mathcal{L}_{\text{LSTM}}\left(\cdot\right)$ in Eq. (30) for LSTMs at the end of training, along with computational times (in seconds) required for training. For the LSTMs, the mean and 10–90% error percentiles (in parentheses) are reported across the 50 training processes with different parameter initializations. Additionally, the MSE loss function of MVARs in Eq. (34) is reported for comparison to the MSE-based loss $\mathcal{L}_{\text{LSTM}}(\cdot)$.
@@
MVAR(4) $1.91\times 10^{-4}$ $1.41\times 10^{-13}$ - $0.0021$
MVAR(9) $1.34\times 10^{-4}$ $7.05\times 10^{-14}$ - $0.0045$
LSTM(4) - - $2.35\ (2.22,2.50)\times 10^{-5}$ 252(236,302)252\ (236,302)
LSTM(9) - - $1.81\ (1.68,1.93)\times 10^{-5}$ 310(302,326)310\ (302,326)
Since the loss function of MVARs is $L_{2}$-based, we additionally compute the MSE in Eq. (34) to enable direct comparison with the MSE-based LSTM loss. MVARs are able to reach much smaller loss function values than LSTMs, providing a hint for their better accuracy. As expected, LSTMs require significantly higher computational time for training in comparison to MVARs. Table 1 further shows that the training of both types of ROMs converges to lower loss function values with increased lag window, albeit requiring significantly more computational time.
One-Step/Open-loop Predictions
Having constructed and trained all components of our framework, we now evaluate its end-to-end performance in one-step predictions in the ambient–density profile–space using the open-loop time-stepper in Eq. (35); note that for the LSTMs, we consider the model with the best training performance among the 50 model runs. We quantify the prediction accuracy via the relative $L_{1}$, $L_{2}$, and $L_{\infty}$ errors, $e^{1,(c)}_{k},\ e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in Eq. (37), computed between the “ground-truth” density profiles $x^{(c)}(t_{k})$ and the predicted/reconstructed ones $\hat{x}^{(c)}(t_{k})$ for the $C$, unseen in training, cases (of different initial conditions) included in the testing set $X_{ts}$ across time steps $k$. Table 2 provides a summary of the relative reconstructed errors in the ambient space, for one-step predictions, for all four ROMs, including the mean error and the 10–90% error percentiles, aggregated over all cases and time steps in the testing set.
Table 2: Open-loop (one-step) prediction errors in the ambient–density profile–space, for the testing set using the trained MVAR and LSTM (best out of 50 training runs) models. The relative reconstructed $L_{1}$, $L_{2}$ and $L_{\infty}$ errors $e^{1,(c)}_{k},\ e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in Eq. (37) are reported for the MVAR(4), MVAR(9), LSTM(4), and LSTM(9) latent dynamics models. Mean and 10–90% error percentiles are shown over all the $C=20$ cases and $K=\{991,996\}$ time steps (for width $w=\{9,4\}$, respectively).
@@
MVAR(4) 0.127(0.072,0.201)0.127\ (0.072,0.201) 0.078(0.045,0.119)0.078\ (0.045,0.119) 0.073(0.039,0.117)0.073\ (0.039,0.117)
MVAR(9) 0.127(0.072,0.200)0.127\ (0.072,0.200) 0.078(0.045,0.118)0.078\ (0.045,0.118) 0.072(0.039,0.116)0.072\ (0.039,0.116)
LSTM(4) 0.280(0.265,0.325)0.280\ (0.265,0.325) 0.167(0.158,0.190)0.167\ (0.158,0.190) 0.138(0.121,0.167)0.138\ (0.121,0.167)
LSTM(9) 0.309(0.297,0.351)0.309\ (0.297,0.351) 0.185(0.174,0.207)0.185\ (0.174,0.207) 0.135(0.115,0.167)0.135\ (0.115,0.167)
As shown in Table 2, the open-loop (one-step) prediction accuracy with the use of MVAR models outperforms that of LSTMs; all LSTM relative errors are approximately twice as high as those of the MVAR models. However, LSTMs demonstrate narrower 10–90% error percentiles in comparison to MVARs, indicating less uncertainty in one-step predictions. Table 2 additionally shows that the accuracy of MVARs does not improve with a longer lag, while that of LSTMs slightly degrades.
(a) Relative Errors with MVAR(4)
(b) Relative Errors with MVAR(9)
(c) Relative Errors with LSTM(4)
(d) Relative Errors with LSTM(9)
Figure 4: One-step (open-loop) prediction errors, in the ambient–density profile–space, for the testing set over time, using trained MVAR and LSTM models. The relative reconstructed $L_{2}$ (red) and $L_{\infty}$ (green) errors $e^{2,(c)}_{k}$ and $e^{\infty,(c)}_{k}$ Eq. (37) (using the one-step prediction in Eq. (35)) are shown for the MVAR(4) panel (a), MVAR(9) panel (b), LSTM(4) panel (c), and LSTM(9) panel (d) latent dynamics models. Mean relative errors (solid) and 10–90% error percentiles (dashed) are shown over $C=20$ cases per time step. For the LSTM models, results correspond to the model achieving the best training performance out of the 50 independent runs.
Fig. 4 displays the average relative reconstructed errors $e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ and their 10–90% percentile ranges, for one-step predictions, over the $C=20$ cases of different, unseen initial conditions considered, across the time steps $k=1,\ldots,1000$. Fig. 4(a,b) depict the results obtained with the MVAR(4) and MVAR(9) models, respectively, both demonstrating a remarkable approximation accuracy. The mean $L_{2}$ reconstructed relative errors are initially $\sim$15% and reduce to $\sim$7% after the first 50s50\ $\text${s} of the simulation. This level of errors matches the infimum approximation error baseline of POD reconstruction, shown in Fig. 3(b). This very high approximation accuracy is not met by the LSTM(4) and LSTM(9) models, shown in Fig. 4(c,d). In particular, the mean relative reconstructed $L_{2}$ errors of the LSTM(4) and LSTM(9) models are $\sim$17% and $\sim$18%, respectively, towards the end of the simulations. Fig. 4 further confirms that the lag window does not affect the one-step prediction accuracy of MVARs, while it slightly reduces that of LSTMs.
Recursive/Closed-loop Predictions
For obtaining long-term (closed-loop) predictions with the proposed framework, we next adopt the recursive closed-loop time-stepper in Eq. (36), which is only supplied with the first $w$ observations and predicts the next time steps in an autoregressive fashion. To quantify the prediction accuracy, we again use the relative reconstructed $L_{1}$, $L_{2}$, and $L_{\infty}$ errors in the ambient space, computed for the cases included in the testing set $X_{ts}$ across time steps. Table 3 provides a summary of these reconstructed errors, including the mean and the 10–90% error percentiles, aggregated over all cases and time steps.
Table 3: Closed-loop/recursive simulation errors in the ambient–density profile–space, for the testing set using the trained MVAR and LSTM (best out of 50 training runs) models. The relative reconstructed $L_{1}$, $L_{2}$ and $L_{\infty}$ errors $e^{1,(c)}_{k},\ e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in Eq. (37) are reported for the MVAR(4), MVAR(9), LSTM(4), and LSTM(9) latent dynamics models. Mean and 10–90% error percentiles are shown over all the $C=20$ cases and $K=\{991,996\}$ time steps (for width $w=\{9,4\}$, respectively).
MVAR(4) 0.231(0.139,0.376)0.231\ (0.139,0.376) 0.153(0.087,0.267)0.153\ (0.087,0.267) 0.141(0.074,0.247)0.141\ (0.074,0.247)
MVAR(9) 0.184(0.114,0.292)0.184\ (0.114,0.292) 0.124(0.073,0.209)0.124\ (0.073,0.209) 0.119(0.062,0.210)0.119\ (0.062,0.210)
LSTM(4) 0.419(0.221,0.648)0.419\ (0.221,0.648) 0.284(0.139,0.512)0.284\ (0.139,0.512) 0.281(0.115,0.501)0.281\ (0.115,0.501)
LSTM(9) 0.433(0.230,0.590)0.433\ (0.230,0.590) 0.310(0.145,0.460)0.310\ (0.145,0.460) 0.285(0.121,0.441)0.285\ (0.121,0.441)
As shown in Table 3, the long-term (closed-loop) prediction accuracy of the proposed framework with MVAR models outperforms significantly that of LSTM models. In addition, MVARs demonstrate narrower 10–90% error percentiles than LSTMs, indicating more reliable long-term predictions. Table 3 further shows that the consideration of a longer lag makes MVARs more accurate, while slightly degrading the accuracy of LSTMs.
@@
We next assessed the evolution/accumulation of the relative reconstructed errors $L_{2}$ and $L_{\infty}$ over time for the closed-loop setting. Fig. 5 depicts the average reconstructed relative errors in the ambient–density profile–space, and their 10–90% error percentile ranges across the multiple cases of different, unseen initial conditions in the testing set for all four ROMs.
@@
Figure 5: Closed-loop/recursive simulation errors in the ambient–density profile–space, across the testing set over time, using the trained MVAR and LSTM models. Panels (a)–(d) display the relative $L_{2}$ (red) and $L_{\infty}$ (green) errors $e^{2,(c)}_{k}$ and $e^{\infty,(c)}_{k}$ in Eq. (37) for MVAR(4), MVAR(9), LSTM(4) and LSTM(9) models, respectively. Panels (e) and (f) provide a system’s “baseline” relative error estimation, obtained from the SFM simulations using $\pm 1\%$ perturbations in the initial conditions for the density profiles (see in Section 3.1). Mean relative error (solid) and 10–90% percentiles (dashed) are shown over $C=20$ cases per time step. For the LSTM models, results correspond to the model achieving the best training performance out of the 50 independent runs.
@@
Indicatively, for our illustrations, to provide a comparative “baseline” of the reconstructed error for tiny perturbations of the density profiles, we quantified the sensitivity of the SFM simulations per se with respect to small perturbations of the initial conditions (see in Section 3.1). Specifically, we introduced $\pm 1\%$ perturbations in the initial conditions of the density profiles of the reference test cases and computed the relative reconstructed errors between these perturbed initial conditions and their unperturbed counterparts. The MVAR models Fig. 5(a,b) demonstrate a remarkable approximation accuracy, with the mean relative $L_{2}$ error remaining bounded below $\sim$14%(17%) for the MVAR(9)(MVAR(4)) models, despite the long-term horizon. The respective 10–90% error percentiles are bounded in (10, 30)% relative errors for the MVAR(4) model, while for the MVAR(9) model, the 10–90% error percentiles reduce further to (7, 21)% relative errors. The mean values of errors, and especially of the MVAR(9) model, is well within the 10–90% error percentiles of the system’s deviation/“baseline” 10–90% $L_{2}$ error percentiles (2, 18)% relative error, as quantified by the SFM simulations of the testing set using the $\pm$ 1% perturbed initial conditions for the density profiles, shown in Fig. 5(e,f). In contrast, the LSTM models Fig. 5(c,d) provide less accurate predictions than the MVAR models, with mean relative $L_{2}$ error plateauing at $\sim$35% for both LSTM(4) and LSTM(9) models, towards the end of the simulation, well above the expected “baseline” error bounds.
For completeness, we directly compare the ground truth normalized density fields $x^{(c)}(t_{k})$ and the predicted fields $\hat{x}^{(c)}(t_{k})$ generated by the closed-loop/recursive simulations. This analysis uses a case from the test dataset featuring a Gaussian initialization corresponding to the 6th row of Table B.2. Figure 6 displays such comparison for the MVAR(9) model at four selected time steps (k=250, 500, 780, 930k=250,\ 500,\ 780,\ 930), during which the pedestrian crowd does not pass through the corridor boundary. Panels (a,c,e,g) show the ground truth at these time steps, while panels (b,d,f,h) present the corresponding MVAR(9) model predictions; analogous results for the MVAR(4), LSTM(4) and LSTM(9) models are provided in Appendix E, Figs. E.1, E.2 and E.3, respectively.
(a) Ground truth density at $t_{250}=62.5\,\text{s}$.
(b) Predicted density at $t_{250}=62.5\,\text{s}$.
(c) Ground truth density at $t_{500}=125\,\text{s}$.
(d) Predicted density at $t_{500}=125\,\text{s}$.
(e) Ground truth density at $t_{780}=195\,\text{s}$.
(f) Predicted density at $t_{780}=195\,\text{s}$.
(g) Ground truth density at time $t_{930}=232.5\,\text{s}$.
(h) Predicted density at $t_{930}=232.5\,\text{s}$.
Figure 6: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ normalized densities via the closed-loop time-stepper in Eq. (36) using the MVAR(9) model for an unseen case of the testing set; initialization at the 6th row of Table B.2. Panels (a), (c), (e), and (g) show the “ground truth” density distribution in $\Omega$, while panels (b), (d), (f), and (h), show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.
@@
The predictions from all four surrogate models successfully preserve the global structure and propagation direction of the ground truth density, demonstrating the framework’s capacity to capture dominant crowd dynamics and produce meaningful long-horizon predictions. The models also correctly respect the obstacle regions. The primary observed discrepancy is a minor underestimation of the peak density values coupled with the introduction of artificial density in low-density regions, a likely consequence of enforcing mass conservation via the POD restriction operator. This artifact is faintest in the MVAR(9) model, and remains only slightly more pronounced in the MVAR(4). By contrast, it becomes considerably stronger in the less accurate LSTM(4) and LSTM(9) models.
MVAR vs. LSTM performance

### v3

@@ 4 Numerical results
Here, we present the results of the proposed framework on predicting pedestrian flow in a rectangular corridor with an obstacle for the unidirectional flow and counterflow configurations described in Section 3.1. We end the section with a comparative discussion about the use of MVAR and LSTM models.
4.1 Unidirectional flow case study
We first considered the unidirectional flow of pedestrians moving from the left to the right of the corridor as shown in Fig. 2(a). As already discussed in Section 3.2, we first collected pedestrian trajectory data using the SFM microscopic agent-based simulator under a wide range of initial conditions. The resulting training dataset $\mathcal{X}_{tr}$ is used to train our framework, while the testing dataset $\mathcal{X}_{ts}$ is used to assess its efficiency and prediction accuracy.
In the first step, we extracted the normalized density fields data $X_{tr},X_{ts}$ from the discrete pedestrian positions in $\mathcal{X}_{tr},\mathcal{X}_{ts}$ using KDE, as described in Section 3.3. Next, we employed POD on $\mathcal{X}_{tr}$ to discover the latent space dimension; see Section 3.4. Following the criterion in Eq. (48), the minimum latent dimension required for retaining more than 99% of the energy in the training dataset is $d=6$ modes, as shown in Fig. 3(a).
@@ 4.1 Unidirectional flow case study
Figure 3: Accuracy of the restriction and lifting operators in Eqs. (34) and (35) for the unidirectional flow case. Panel (a) shows the minimum number of POD modes (colored diamonds) required for retaining the desired percentage of energy in the training data set. Panel (b) depicts the average (over the $C=20$ cases with different initial conditions) reconstruction error $e^{2,rec}_{k}$ in Eq. (49) with $d=6$ POD modes, in time for the training and testing datasets.
Using the discovered POD basis $U_{d}$, we next constructed the restriction and lifting operators in Eqs. (34) and (35). To assess their accuracy, we computed the relative $L_{2}$ reconstruction error $e^{2,rec}_{k}$ in Eq. (49) across the $k=1,\ldots,K$ time steps of each trajectory. Fig. 3(b) displays the average relative error over all the $C=20$ cases with different initial conditions included in either the training or testing datasets. While initially the relative reconstruction error is $\sim$21%, it decreases to 4% at the first 20s20\ $\text${s} and further reduces to $\sim$3% after 50s50\ $\text${s} for both seen (training set) and unseen (testing set) data, indicating high accuracy of the restriction and lifting operators with a significantly reduced latent space dimension $d=6$. We note that this level of reconstruction error serves as an infimum approximation error baseline, i.e., the best possible error for the proposed framework at the ambient–density profiles–space. ROM forecasts in the long term are therefore expected to surpass this baseline error due to the additional error contributions from model-form error and training/generalization errors.
We next trained MVAR and LSTM models to learn ROMs for the latent dynamics (see Section 2.2.3), where the latent variables $\boldsymbol{y}_{c}(t_{k})\in\mathbb{R}^{6}$ of the $c$ case at timestep $t_{k}$ are obtained from the density profiles of the training set via the restriction operator in Eq. (34). To determine the lag window, we employed the BIC and AIC criteria described in D; BIC indicated a lag window of $w=4$, whereas AIC favored a wider (as expected) $w=9$. Both windows were used for training both MVAR and LSTM models, denoted from now on as MVAR(4), MVAR(9), LSTM(4), and LSTM(9) models. The training data sets were divided into features $Y_{f}=\{\boldsymbol{y}_{c}(t_{k-w}),\ldots,\boldsymbol{y}_{c}(t_{k-1})\}$ and targets $Y_{t}=\{\boldsymbol{y}_{c}(t_{k})\}$.
@@
Table 1: Training results of MVAR(4), MVAR(9), LSTM(4) and LSTM(9) models at the latent space for the unidirectional flow case. Loss functions $\mathcal{L}_{\text{MVAR}}\left(\cdot\right)$ in Eq. (41) for MVARs and $\mathcal{L}_{\text{LSTM}}\left(\cdot\right)$ in Eq. (46) for LSTMs are reported at the end of training, along with computational times (in seconds) required for training. For the LSTMs, the mean and 10–90% error percentiles (in parentheses) are reported across the 50 training processes with different parameter initializations. Additionally, the MSE loss function of MVARs in Eq. (50) is reported for comparison to the MSE-based loss $\mathcal{L}_{\text{LSTM}}(\cdot)$.
@@
MVAR(4) $4.65\times 10^{-8}$ $3.89\times 10^{-13}$ - $0.0013$
MVAR(9) $2.04\times 10^{-8}$ $1.72\times 10^{-13}$ - $0.0017$
LSTM(4) - - $7.12\ (6.27,8.16)\times 10^{-9}$ 62(55,66)62\ (55,66)
LSTM(9) - - $1.51\ (1.30,1.81)\times 10^{-8}$ 86(77,89)86\ (77,89)
Since the loss function of MVARs is $L_{2}$-based, we additionally compute the MSE in Eq. (50) to enable direct comparison with the MSE-based LSTM loss. MVARs converge to much smaller loss function values in comparison to LSTMs. As expected, LSTMs require significantly higher computational time for training in comparison to MVARs. Table 1 further shows that the training of both types of ROMs with increased lag window requires more computational time.
Recursive/Closed-loop
Predictions Having constructed and trained all components of our framework, we evaluated its end-to-end performance in the ambient–density profile–space. While the ROMs were trained in an open-loop setting, we assess them in the practical closed-loop setting (see Eq. (51)), where only the first $w$ observations are supplied and all subsequent predictions are generated autoregressively. For LSTM models, we consider the best-performing model from the 50 training runs. We quantify the prediction accuracy via the relative $L_{1}$, $L_{2}$, and $L_{\infty}$ errors, $e^{1,(c)}_{k},\ e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in Eq. (52), computed between the “ground-truth” density profiles $x^{(c)}(t_{k})$ and the predicted/reconstructed ones $\hat{x}^{(c)}(t_{k})$ for the $C$, unseen in training, cases (of different initial conditions) included in the testing set $X_{ts}$ across time steps $k$.
Table 2 provides a summary of the closed-loop reconstructed errors, including the mean and the 10–90% error percentiles, aggregated over all cases and time steps.
Table 2: Closed-loop/recursive prediction relative errors in the ambient–density profile–space, for the testing set using the trained MVAR and LSTM (best out of 50 training runs) models for the unidirectional flow case. The relative reconstructed $L_{1}$, $L_{2}$ and $L_{\infty}$ errors $e^{1,(c)}_{k},\ e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in Eq. (52) are reported for the MVAR(4), MVAR(9), LSTM(4), and LSTM(9) latent dynamics models. Mean and 10–90% error percentiles are shown over all the $C=20$ cases and $K=\{991,996\}$ time steps (for width $w=\{9,4\}$, respectively).
@@
MVAR(4) 0.180(0.085,0.246)0.180\ (0.085,0.246) 0.153(0.077,0.211)0.153\ (0.077,0.211) 0.172(0.095,0.242)0.172\ (0.095,0.242)
MVAR(9) 0.163(0.087,0.228)0.163\ (0.087,0.228) 0.140(0.079,0.194)0.140\ (0.079,0.194) 0.160(0.092,0.230)0.160\ (0.092,0.230)
LSTM(4) 0.177(0.128,0.235)0.177\ (0.128,0.235) 0.156(0.115,0.209)0.156\ (0.115,0.209) 0.185(0.133,0.243)0.185\ (0.133,0.243)
LSTM(9) 0.165(0.112,0.223)0.165\ (0.112,0.223) 0.142(0.110,0.190)0.142\ (0.110,0.190) 0.166(0.116,0.227)0.166\ (0.116,0.227)
As shown in Table 2, the proposed framework with the use of MVAR models slightly outperforms that of LSTMs in the closed-loop setting across error metrics. The MVAR(9) achieves the lowest errors, indicating accuracy in long-term predictions. The MVAR(4) also performs well, with broader percentiles across test cases. Increasing the lag window from $w=4$ to $w=9$ slightly improves MVAR accuracy, demonstrating the benefit of incorporating longer temporal dependencies within the linear autoregressive structure. The LSTM models exhibit slightly higher errors than MVARs, with tighter 10–90% percentiles, while the increase of lag window shows a similar to MVARs improvement.
@@
We next assessed the evolution/accumulation of the relative reconstructed errors $L_{2}$ and $L_{\infty}$ over time. Figure 4 depicts the average reconstructed relative errors in the ambient–density profile–space, and their 10–90% error percentile ranges across the multiple cases of different, unseen initial conditions in the testing set for all four ROMs.
@@
Figure 4: Closed-loop/recursive simulation errors in the ambient–density profile–space, across the testing set over time, using the trained MVAR and LSTM models for the unidirectional flow case. Panels (a)–(d) display the relative $L_{2}$ (cyan) and $L_{\infty}$ (green) errors $e^{2,(c)}_{k}$ and $e^{\infty,(c)}_{k}$ in Eq. (52) for MVAR(4), MVAR(9), LSTM(4) and LSTM(9) models, respectively. Panels (e) and (f) provide a system’s “baseline” relative error estimation, obtained from the SFM simulations using $\pm 1\%$ perturbations in the initial conditions for the density profiles (see in Section 3.1). Mean relative error (solid) and 10–90% percentiles (dashed) are shown over $C=20$ cases per time step. For the LSTM models, results correspond to the model achieving the best training performance out of the 50 independent runs.
@@
Indicatively, for our illustrations, to provide a comparative “baseline” of the reconstructed error for tiny perturbations of the density profiles, we quantified the sensitivity of the SFM simulations per se with respect to small perturbations of the initial conditions (see in Section 3.1). Specifically, we introduced $\pm 1\%$ perturbations in the initial conditions of the density profiles of the reference test cases and computed the relative reconstructed errors between these perturbed initial conditions and their unperturbed counterparts, as shown in Fig. 4(e,f).
The time evolution in Fig. 4 reveals a characteristic pattern across all models: errors are higher initially, decrease drastically during the first $20\,\text{s}$ as the POD reconstruction error baseline diminishes, and then slowly increase towards the end of the simulation due to error accumulation in the recursive forecasting. The MVAR models in Fig. 4(a,b) demonstrate remarkable approximation accuracy with the mean relative $L_{2}$ error remaining bounded below $\sim$14%(15%) for the MVAR(9)(MVAR(4)) models by the end of the simulation horizon. The respective 10–90% error percentiles are bounded in (7, 25)% errors for the MVAR(4) model, while for the MVAR(9) model they tighten further to (10, 20)%. Both MVAR models achieve errors that compare favorably with the intrinsic system baseline obtained from $\pm 1\%$ perturbed SFM simulations, which shows mean $L_{2}$ error of $\sim 8\%$ and 10–90% percentiles $(3,32)\%$ (Fig. 4(e,f)). The LSTM models in Fig. 4(c,d) provide slightly less accurate predictions than the MVAR models. By the end of the simulation, the LSTM(4) mean relative $L_{2}$ error reaches approximately $16\%$, with 10–90% percentiles widening to $(11,25)\%$. The LSTM(9) improves marginally, with mean error around $13\%$ and tighter percentiles $(10,18)\%$. The MVAR and LSTM models demonstrate similar trends, with MVARs providing slightly higher accuracy; a result confirming the averaged in time results reported in Table 2.
@@
For completeness, we directly compared the ground truth normalized density fields $x^{(c)}(t_{k})$ and the predicted fields $\hat{x}^{(c)}(t_{k})$ generated by the closed-loop/recursive simulations. The predictions of the proposed framework for all ROMs (see Figs E.2E.5) successfully preserve the global structure and propagation direction of the ground truth density, demonstrating the framework’s capacity to capture dominant crowd dynamics and produce meaningful long-horizon predictions. The spatial distribution of the absolute errors between ground truth and closed-loop predictions (see Fig. E.6) further shows that the all models correctly respect the obstacle regions. The primary observed discrepancy is a minor underestimation of the peak density values, and of density in regions immediately downstream and upstream of these peaks coupled with the introduction of artificial density in low-density regions; the latter is a likely consequence of enforcing mass conservation via the POD restriction operator. These artifacts are fainter in the MVAR(9) and LSTM(9) models, and become slightly stronger in the less accurate MVAR(4) and LSTM(4) models, showing a lessening effect as time passes (see Fig. E.6).
Finally, the open-loop (one-step) prediction results are provided in E; see Table E.1 and Fig. E.1. Similarly to the closed-loop findings above, all four ROMs achieve comparable accuracy in the open-loop setting, with mean $L_{2}$ errors of approximately $0.038$ across models and narrow 10–90% percentiles. This indicates that while both models capture the local dynamics very accurately in one-step predictions ($L_{2}$ errors just above the POD baseline), their long-term forecasting is, in practice, affected by error accumulation that can degrade their performance.
4.2 Counterflow case study
For the second case study, we applied the proposed framework in a counterflow configuration in which two interacting populations are divided equally into two groups with opposing flow directions; see Fig. 2(b). Following the same procedure as in the unidirectional flow case (for implementation details see Sections 3.23.4), we constructed the restriction and lifting operators in Eqs. (36) and (37) by retaining $d_{1}=6$ and $d_{2}=8$ POD modes per group $l=1,2$ and $m=4$ cross-covariance modes. The accuracy of the resulting operators is very high, with relative $L_{2}$ errors reducing to $3-4\%$ after the first $50\,\text{s}$ for both groups (see Fig. F.1 in F).
With latent variables $\boldsymbol{y}_{c}(t_{k})=[\boldsymbol{y}^{(1)}_{c}(t_{k}),\boldsymbol{y}^{(2)}_{c}(t_{k})]\in\mathbb{R}^{24}$ determined, we trained MVAR and LSTM models following Section 3.5. The AIC-identified lag was $w=10$, and stationarity was verified (ADF test). The MVAR(10) again converged to lower loss function values with substantially less computational time than the LSTM(10) one (see Table F.1 in F), though both models show deteriorated performance compared to the unidirectional flow case.
To evaluate the end-to-end performance of the proposed framework in the ambient–density profile–space in long-term (closed-loop) predictions, we followed the same quantification procedure as in Section 4.1. We again observe that the closed-loop prediction accuracy with the use of MVAR is superior to that of LSTM (see Table F.2 in F). In particular, the mean, overall testing cases and time steps, reconstructed $L_{2}$ error of the MVAR(10) model is $8\%$ with tight 10–90% percentiles, while that of the LSTM(10) model is $\sim 10\%$ with wider percentiles for both population groups. The detailed error evolution over time is displayed in Fig. 5. The MVAR(10) model shows mean $L_{2}$ errors dropping to $\sim 9-10\%$ within the first $20\,\text{s}$ for both population groups. These errors remain bounded over the long-time horizon at values slightly over $10\%$, exhibiting high accuracy in comparison to the baseline accuracy of pure restriction and lifting operators ($3-4\%$ after the first $50\,\text{s}$ for both groups; see Fig. F.1). In contrast, LSTM(10) exhibits slightly larger errors that stabilize at $13-14\%$ by the simulation end. Across both models, group 2 shows slightly higher errors and wider uncertainty bands.
(a) Relative Errors with MVAR(10) for group 1
(b) Relative Errors with MVAR(10) for group 2
(c) Relative Errors with LSTM(10) for group 1
(d) Relative Errors with LSTM(10) for group 2
Figure 5: Closed-loop/recursive simulation errors in the ambient–density profile–space, across the testing set over time, using the MVAR and LSTM models for the counterflow case. Panels (a,b) and (c,d) display the relative reconstructed $L_{2}$ (cyan) and $L_{\infty}$ (green) errors $e^{2,(c)}_{k}$ and $e^{\infty,(c)}_{k}$ in Eq. (52) for MVAR(10) and LSTM(10) models, respectively. The left and right column panels show errors for group 1 and 2, respectively. Mean relative errors (solid) and 10–90% error percentiles (dashed) are reported over $C=20$ cases per time step. For the LSTM model, results correspond to the run achieving the best training performance out of the 50 independent initializations.
Additionally, we provide a direct comparison of the ground truth densities with their closed-loop predictions for both groups (see Figs. F.3 and F.4 in F). Both MVAR and LSTM models accurately reproduce the two dominant density lobes and capture migration and alignment patterns of the two interacting groups. These results, together with the spatial distribution of the absolute errors between ground truth and closed-loop predictions (see Fig. F.5), further show that both models respect obstacle regions, without significant errors accumulating in the immediate obstacle proximity, while accurately reproducing localized congestion zones formed upstream and downstream of the obstacle. Minor discrepancies include peak attenuation and spatial dispersion of the Gaussian-like lobe structure of each group, which are slightly more pronounced with the LSTM(10) model, especially as time progresses; a finding that is consistent with LSTM’s higher error metrics compared to MVAR in Fig. 5.
For completeness, we also provide the open-loop/one-step prediction results in Table F.2 and Fig. F.2 of F. As in the unidirectional flow case, both MVAR and LSTM models achieve comparable accuracy with mean $L_{2}$ errors $\sim 4-6\%$ and very narrow 10–90% percentiles. This very accurate performance in one-step predictions demonstrates once again the effect of error accumulation introduced in the recursive predictions for long-term horizons.
Finally, the proposed framework achieves higher accuracy in the counterflow case than in the unidirectional flow case, and this difference is consistent across ROM model selections. This may be attributed to the larger latent dimension ($24$ vs $6$), which provides a richer representation space. However, this comes at increased computational cost, highlighting the accuracy-efficiency trade-off inherent in dimensionality selection.
4.3 MVAR vs. LSTM performance